What is the uninflated present worth of $5000 in 3 years if the average inflation rate is 6% and i is 10%?
Answer #1

Inflation adjusted interest rate, = 0.06 +0.10 + (0.06)(0.12 = 0.166 P= ($5000) (P/F,16.6%,3) = ($5000) (4 + din = ( $5000)